While running, a 74 kg student generates thermal energy at a rate of 1300 W. To maintain a constant body temperature of 37.0 ∘C, this energy must be removed by perspiration or other mechanisms.

Part A If these mechanisms failed and the heat could not flow out of the student's body, for what amount of time could a student run before irreversible body damage occurred? (Protein structures in the body are damaged irreversibly if the body temperature rises to 44.0 ∘C or above. The specific heat capacity of a typical human body is 3480 J/(kg⋅K), slightly less than that of water. The difference is due to the presence of protein, fat, and minerals, which have lower specific heats.)
Answer #1

Solution : amount of heat required to increase body temperature from 37 to 44 degree Celsius.

Q = Ms dT = 74 x 3480 x(44-37) = 1802640 Joules

Now energy generated per unit time = 1300 Joules.

Time = total energy/ power = 1802640 / 1300 = 1386.65 seconds